# Learning from Heterogeneity: A Dynamic Learning Framework for Hypergraphs

### v2

@@ Abstract
Index Terms:
Heterogeneous Hypergraph, Representation Learning, Hypergraph Generation, Classification
I Introduction
@@
The remainder of this paper is organized as follows. Section II reviews the related works of graph learning on homogenous, heterogenous, and hypergraphs. Section III provides the preliminary knowledge of hypergraphs, while Section IV introduces the detail of our proposed LFH framework, along with analyzes and discussions. The experimental results are reported and analyzed in Section V. Finally, Section VI concludes the paper.
II Related Work
II-A Homogeneous Graph Learning
@@
II-B Heterogeneous Graph Learning
@@ II-C Hypergraph Learning
III Preliminaries and problem formulation
In this section, we first introduce the relevant definitions and then formulate the key problems associated with representation learning on heterogeneous hypergraphs. The notations used in this paper are listed in Table I.
@@ III Preliminaries and problem formulation
TABLE I: Table of key notations.
@@ E Hyperedges embeddding
$K$ The number of attention heads
IV Heterogeneous Hypergraph
@@
IV-A Initial Node Embedding
As demonstrated in Fig. 2, the input of this step is the heterogeneous pairwise graph. It is of great importance to generate high-quality initial node embedding to better exploit the implicit data relations in the hypergraph. Intuitively, fusing pairwise information from a heterogeneous pairwise graph into the initial node embeddings can help capture high-order correlations among the nodes and thus affects the construction of the hypergraph. Given a heterogeneous pairwise graph $\dot{\mathcal{G}}=\{\dot{\mathcal{V}},\dot{\mathcal{E}},\dot{\mathcal{T}_{e}},\dot{\mathcal{T}_{v}}\}$ and its raw node features $\dot{\emph{X}}=\left[\dot{\emph{x}}_{1},\dot{\emph{x}}_{2},...,\dot{\emph{x}}_{N}\right]\in\mathcal{R}^{D\times N}$, where the raw feature of node $v_{i}\in\mathcal{V}$ is $\dot{\emph{x}}_{i}\in\mathcal{R}^{D}$. The pairwise fusion can be implemented by any GNN model that operates on heterogeneous pairwise graphs. Formally, the pairwise fusion process can be defined as:
@@ IV-A Initial Node Embedding
IV-B Heterogeneous Hyperedge Generation
@@
IV-C Dynamic Hypergraph Learning
@@
IV-C1 Hyperedge Embedding Updating
@@
IV-C2 Multi-head Attention Node Embedding Updating
@@
where the dimension of $\emph{X}\left(v^{m}_{i}\right)$ is $\mathcal{R}^{d\times 1}$, and $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}\in\mathcal{R}^{\frac{d}{K}\times d}$ represents the projection matrix on the $h$-th attention head, $h\in\left[1,K\right]$. $K$ is the number of attention heads. The master node is projected to generate the $h$-th query vector $\emph{Q}^{h}\left(v^{m}_{i}\right)$ by the projection matrix $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}$. Note the projection matrices are distinguished on hyperedge types. Similarly, we also project the hyperedge $e^{1}\left(v^{m}_{i}\right)$ through $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ into $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ as:
@@
where $\emph{E}\left(e^{1}\left(v^{m}_{i}\right)\right)\in\mathcal{R}^{d\times 1}$ is the embedding of hyperedge $e^{1}\left(v^{m}_{i}\right)$. The dimension of $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ is $\mathcal{R}^{\frac{d}{K}\times d}$ while the dimension of the output $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ is $\mathcal{R}^{\frac{d}{K}}$.
@@
Eq. 11 derives the attention value of the $h$-th head between $v^{m}_{i}$ and $e^{1}\left(v^{m}_{i}\right)$. The type-specific learnable matrix $\Theta^{att}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}\in\mathcal{R}^{\frac{d}{K}\times\frac{d}{K}}$ for edge type $\psi\left(e^{1}\left(v^{m}_{i}\right)\right)$ represents the learnable semantic information for each edge type. $\mu$ is a scaling factor for different hyperedge types. Moreover, since the magnitude of K and Q can increase the attention value significantly and eventually lead to the gradient explosion problem, we divide the obtained value by $\sqrt{d}$ to stabilize the training process.
@@
$\tilde{\emph{w}}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right)=\mathop{\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}}\limits_{h\in\left[1,K\right]}att^{h}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right)$ (12)
@@
where $\emph{w}\left(v^{m}_{i},e\right)=\left[\emph{w}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right),...,\emph{w}\left(v^{m}_{i},e^{M_{i}}\left(v^{m}_{i}\right)\right)\right]\in\mathcal{R}^{K\times M_{i}}$, with entries are normalized among all hyperedges per head (illustrated in the right part of Fig. 3). The attentive aggregation of head $h$ for node $v^{m}_{i}$ is formulated as:
@@
$w^{h}\left(v^{m}_{i},e^{k}\left(v^{m}_{i}\right)\right)$ is the scalar element of $h$-th head in the weight vector $\emph{w}\left(v^{m}_{i},e^{k}\left(v^{m}_{i}\right)\right)$ of hyeredge $e^{k}\left(v^{m}_{i}\right)$, and the dimension of $\emph{z}_{i}^{h}$ is $\mathcal{R}^{\frac{d}{K}}$. After concatenating all $K$ heads, we obtain the updated node embedding of $v^{m}_{i}$ through a shallow multi-layer perceptron (MLP), which is denoted as:
$\emph{Z}^{\mathcal{V}}\left(v^{m}_{i}\right)=\textit{MLP}\left(\mathop{\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}}\limits_{h\in\left[1,K\right]}\emph{z}^{h}_{i}\right)$ (15)
IV-D Loss Function
@@
where $\mathcal{L}_{recon}$ is defined in Eq. 6. The process of reconstructing the master node using slave nodes can uncover implicit relationships between the master node and others. However, as pointed out above, this process is sensitive to noisy nodes. Over-reliance on reconstruction loss compromises the performance of our model. To enhance the generalization ability, we need to use a weight hyperparameter $\alpha$, to balance the effects of loss of supervised downstream task and reconstruction loss. The full process of the model, as described in Algorithm 1, and the detailed analysis of computational cost are provided in Appendix.
IV-E Discussion
@@ V Experiment
  • RQ3: How much do the node embedding size and the number of multi-head $K$ affect the final results?
@@
V-A Experimental Setup
TABLE II: Statistics of the datasets.
@@ V-A Experimental Setup
V-A1 Dataset
The experiments are conducted on three datasets, which are popularly used in classification tasks. The details of these datasets are summarized in Table II.
@@ V-A1 Dataset
TABLE III: Shared parameter setup.
@@
V-A2 Baseline Models and Configurations
@@
The shared parameter setup of the framework can be found in Table III. We randomly split all datasets into train/validate/test with the ratio of 0.2/0.1/0.7, respectively. It is worth mentioning that the hyperparameters of all baseline models are set to the optimal values reported in their original papers. All models are trained with a fixed 100 epochs, using an early stopping strategy when the performance on the validation set is not improved for 30 consecutive epochs. All trainable parameters of the neural network are initialized through Xavier [45] and optimized using Adam [44] with the learning rate 2e-3. The dropout rate is set to be 0.3 [46]. The optimal selection of hyperparameters of our model is further analysed in the following section.
TABLE IV: Performance of different models on node classification task.
@@
V-B Performance Analysis (RQ1)
We evaluate the performance of our proposed framework on node classification and link prediction tasks.
V-B1 Node Classification (RQ1.1)
In the node classification task, we initially perform a random sampling of nodes at varying percentages (10%, 30%, 50%, 70%) of the nodes from the three datasets (ACM, IMDB and DBLP). Then, any edges connected to nodes that are not sampled are eliminated. Subsequently, any sampled nodes that lack neighboring nodes are also removed. It is observed that a lower data sampling ratio correlates with an increased introduction of noise into the training dataset, due to the loss of original information. The best F1 results among all datasets are demonstrated in Table IV, in which the highest scores are marked in bold. We compared three different categories of graph learning models: homogeneous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraph learning models. Our proposed LFH achieves performance gain over other baseline models in all datasets by 2%-38%. Our model outperforms all baseline models in all datasets by an average improvement of 12.9%. Compared to four homogeneous pairwise GNN baselines GCN, GAT, GraphSAGE and GraphSAINT, the average F1 score improvement of LFH for all datasets is 14.4%. Regarding the cherry-picked heterogeneous pairwise graph learning models and hypergraph learning models, our model achieved an average improvement of 9.7% and 14.6%, respectively, compared to the best-recorded results in all datasets. It is worth noting that HGNN+ [20] is the only hypergraph learning baseline that utilizes pairwise information for hyperedge generation, resulting in better performance compared to other hypergraph learning models, except for our model. Furthermore, the training process of our model, which incorporates type-specific hyperedge generation and attentive embedding updates that exploits the heterogeneity of the graph, enables LFH to improve HGNN+ by an average of 5.7% in all datasets. Additionally, compared to the latest reporting heterogeneous learning model PC-HGN [8], the best results of our model gains an average increase by 5.2%. In general, we observe that our model outperforms all baselines for different data sample ratios (DSR), which is shown in Table IV. When the data sample ratios are set to be 70%, 50%, 30% and 10%, our model outperforms other models by 10.1%, 10.2%, 10.9% and 16.7% on average, respectively. It is worth noting that our model achieves higher average improvements when the data sampling ratio is set to 10%, and then the margin of average improvements declines as the data sampling ratio increases.
TABLE V: Performance of different models on link prediction.

### v3

@@ Abstract
{IEEEImpStatement}
This research proposes a novel hypergraph learning framework, namely Learning from Heterogeneity (LFH), for dynamically constructing the hypergraph while adaptively learning heterogeneous hypergraph representations. LFH generates the high-quality initial node embedding using the designated pairwise fusion function, exploiting pairwise graph information at most. To encode high-order data relations, this research designs a dynamic learning process to generate different types of implicit hyperedges to construct the hyper graph and dynamically adapts the hypergraph construction during the learning process. The framework iteratively updates node embeddings through a type-specific multi-head attention mechanism that adaptively learns the importance of heterogeneous hyperedges during the updates, thereby encoding the heterogeneity of the topology into the embedding space. LFH excels in both node classification and edge prediction tasks compared to state-of-the-art baselines.
{IEEEkeywords}
Classification, Heterogeneous Hypergraph, Hypergraph Generation, Representation Learning
1 Introduction
[…]
@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 reviews the related works of graph learning on homogenous, heterogenous, and hypergraphs. Section 3 provides the preliminary knowledge of hypergraphs, while Section 4 introduces the detail of our proposed LFH framework, along with analyzes and discussions. The experimental results are reported and analyzed in Section 5. Finally, Section 6 concludes the paper.
2 Related Work
2.1 Homogeneous Graph Learning
@@
2.2 Heterogeneous Graph Learning
@@ 2.3 Hypergraph Learning
To elucidate the tasks addressed in relevant works, hypergraph learning methods can be divided into two categories: methods for relational data and non-relational data. The methods for relational data focus on tasks that naturally involve explicit correlations among data, with well-known examples including social network tasks [37, 17, 16] and social network tasks [32, 33, 35]. Previous works usually define handcrafted rules [33, 17, 16] or utilize a similarity function to select neighboring nodes based on existing pairwise relations to construct the hypergraph [32, 35, 37]. On the other hand, methods [34, 19, 38] for non-relational data focus on datasets containing independent data points, such as images, where there are no explicit correlations among different images. These methods utilize unsupervised clustering strategies to construct hypergraphs, from which hypergraph learning models can be applied.
@@
3 Preliminaries and problem formulation
In this section, we first introduce the relevant definitions and then formulate the key problems associated with representation learning on heterogeneous hypergraphs. The notations used in this paper are listed in Table 1.
@@ 3 Preliminaries and problem formulation
Table 1: Table of key notations.
@@ E Hyperedges embeddding
$H$ The number of attention heads
4 Heterogeneous Hypergraph
@@
4.1 Initial Node Embedding
As demonstrated in Fig. 2, the input of this step is the heterogeneous pairwise graph. It is of great importance to generate high-quality initial node embedding to better exploit the implicit data relations in the hypergraph [20, 16]. Intuitively, fusing pairwise information from a heterogeneous pairwise graph into the initial node embeddings can help capture high-order correlations among the nodes and thus affects the construction of the hypergraph. Given a heterogeneous pairwise graph $\dot{\mathcal{G}}=\{\dot{\mathcal{V}},\dot{\mathcal{E}},\dot{\mathcal{T}_{e}},\dot{\mathcal{T}_{v}}\}$ and its raw node features $\dot{\emph{X}}=\left[\dot{\emph{x}}_{1},\dot{\emph{x}}_{2},...,\dot{\emph{x}}_{N}\right]\in\mathcal{R}^{D\times N}$, where the raw feature of node $v_{i}\in\mathcal{V}$ is $\dot{\emph{x}}_{i}\in\mathcal{R}^{D}$. The pairwise fusion can be implemented by any GNN model that operates on heterogeneous pairwise graphs. Formally, the pairwise fusion process can be defined as:
@@ 4.1 Initial Node Embedding
4.2 Heterogeneous Hyperedge Generation
@@
4.3 Dynamic Hypergraph Learning
@@
4.3.1 Hyperedge Embedding Updating
@@
4.3.2 Multi-head Attention Node Embedding Updating
@@
where the dimension of $\emph{X}\left(v^{m}_{i}\right)$ is $\mathcal{R}^{d\times 1}$, and $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}\in\mathcal{R}^{\frac{d}{H}\times d}$ represents the projection matrix on the $h$-th attention head, $h\in\left[1,H\right]$. $H$ is the number of attention heads. The master node is projected to generate the $h$-th query vector $\emph{Q}^{h}\left(v^{m}_{i}\right)$ by the projection matrix $Q\text{-}Proj^{h}_{\phi\left(v^{m}_{i}\right)}$. Note the projection matrices are distinguished on hyperedge types. Similarly, we also project the hyperedge $e^{1}\left(v^{m}_{i}\right)$ through $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ into $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ as:
@@
where $\emph{E}\left(e^{1}\left(v^{m}_{i}\right)\right)\in\mathcal{R}^{d\times 1}$ is the embedding of hyperedge $e^{1}\left(v^{m}_{i}\right)$. The dimension of $K\text{-}Proj^{h}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}$ is $\mathcal{R}^{\frac{d}{H}\times d}$ while the dimension of the output $\emph{K}^{h}\left(e^{1}\left(v^{m}_{i}\right)\right)$ is $\mathcal{R}^{\frac{d}{H}}$.
@@
Eq. 11 derives the attention value of the $h$-th head between $v^{m}_{i}$ and $e^{1}\left(v^{m}_{i}\right)$. The type-specific learnable matrix $\Theta^{att}_{\psi\left(e^{1}\left(v^{m}_{i}\right)\right)}\in\mathcal{R}^{\frac{d}{H}\times\frac{d}{H}}$ for edge type $\psi\left(e^{1}\left(v^{m}_{i}\right)\right)$ represents the learnable semantic information for each edge type. $\mu$ is a scaling factor for different hyperedge types. Moreover, since the magnitude of K and Q can increase the attention value significantly and eventually lead to the gradient explosion problem, we divide the obtained value by $\sqrt{d}$ to stabilize the training process.
@@
$\tilde{\emph{w}}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right)=\mathop{\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}}\limits_{h\in\left[1,H\right]}att^{h}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right)$ (12)
@@
where $\emph{w}\left(v^{m}_{i},e\right)=\left[\emph{w}\left(v^{m}_{i},e^{1}\left(v^{m}_{i}\right)\right),...,\emph{w}\left(v^{m}_{i},e^{M_{i}}\left(v^{m}_{i}\right)\right)\right]\in\mathcal{R}^{H\times M_{i}}$, with entries are normalized among all hyperedges per head (illustrated in the right part of Fig. 3). The attentive aggregation of head $h$ for node $v^{m}_{i}$ is formulated as:
@@
$w^{h}\left(v^{m}_{i},e^{k}\left(v^{m}_{i}\right)\right)$ is the scalar element of $h$-th head in the weight vector $\emph{w}\left(v^{m}_{i},e^{k}\left(v^{m}_{i}\right)\right)$ of hyeredge $e^{k}\left(v^{m}_{i}\right)$, and the dimension of $\emph{z}_{i}^{h}$ is $\mathcal{R}^{\frac{d}{H}}$. After concatenating all $H$ heads, we obtain the updated node embedding of $v^{m}_{i}$ through a shallow multi-layer perceptron (MLP), which is denoted as:
$\emph{Z}^{\mathcal{V}}\left(v^{m}_{i}\right)=\textit{MLP}\left(\mathop{\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}}\limits_{h\in\left[1,H\right]}\emph{z}^{h}_{i}\right)$ (15)
4.4 Loss Function
@@
where $\mathcal{L}_{recon}$ is defined in Eq. 6. The process of reconstructing the master node using slave nodes can uncover implicit relationships between the master node and others. However, as pointed out above, this process is sensitive to noisy nodes. Over-reliance on reconstruction loss compromises the performance of our model. To enhance the generalization ability, we need to use a weight hyperparameter $\alpha$, to balance the effects of loss of supervised downstream task and reconstruction loss. The full process of LFH is described in Algorithm 1.
Algorithm 1 Pseudocode of LFH training process
0:  The heterogeneous pairwise graph $\dot{\mathcal{G}}$; the raw node feature $\dot{\emph{X}}$; the set of node $\mathcal{V}$; the set of node type $\mathcal{T}_{v}$; the number of attention heads $H$; the weight hyperparameter $\alpha$
0:  Embedding of all nodes $\emph{Z}^{\mathcal{V}}$
1:  X $\leftarrow$ Generate initial node embeddings by pairwise fusion (Eq.3)
2:  for each master node $v^{m}_{i}\in\mathcal{V}$ do
3:     for each node type $k\in\mathcal{T}_{v}$ do
4:        $\emph{X}(v^{m}_{i})$leftarrow The embedding of the master node $v^{m}_{i}$
5:        $\hat{\emph{X}_{k}}(v^{m}_{i})$leftarrow Reconstruct $\emph{X}(v^{m}_{i})$ by the slave node set of the node type $k$ (Eq.4)
6:        Compute the reconstruction error between $\emph{X}(v^{m}_{i})$ and $\hat{\emph{X}_{k}}(v^{m}_{i})$ through Eq.5
7:     end for
8:  end for
9:  $\mathcal{L}_{recons}$leftarrow Calculate the loss of reconstruction by Eq.6
10:  H $\leftarrow$ Generate the incidence matrix of hypergraph by Eq.7
11:  for each master node $v^{m}_{i}\in\mathcal{V}$ do
12:     $\mathcal{E}(v^{m}_{i})$leftarrow Hyperedges embedding associated with $v^{m}_{i}$ updated by Eq.8
13:  end for
14:  for each master node $v^{m}_{i}\in\mathcal{V}$ do
15:     for each $e(v^{m}_{i})\in\mathcal{E}(v^{m}_{i})$ do
16:        for each $h=1,...,H$ do
17:           $w^{h}(v^{m}_{i},e(v^{m}_{i}))$leftarrow Compute the normalized multi-head attention as the weight of the $h$-th head between $v^{m}_{i}$ and $e(v^{m}_{i})$ by Eq.9 to Eq.13
18:        end for
19:     end for
20:  end for
21:  $\emph{Z}^{\mathcal{V}}$leftarrow Update node embedding from Eq.14 to Eq.15
22:  $\mathcal{L}_{label}$leftarrow Compute the loss of supervised downstream task by Eq.16.
23:  Compute the united loss $\mathcal{L}_{u}=(1-\alpha)\times\mathcal{L}_{label}+\alpha\times\mathcal{L}_{recons}$ (Eq.17), and update all trainable weights.
4.5 Model Analysis and Discussion
We analyze the computational cost and summarize the advantages of LFH as below:
4.5.1 Computational Cost
In the process of heterogeneous hypergraph generation, the computational cost mainly lies in the reconstruction of the master node from its candidate slave node set (Eq. 4). As it is essential to generate hyperedges of different types for each node, and the computational cost of heterogeneous hypergraph generation is $\emph{O}\left(d|\mathcal{T}_{v}|N\right)$, where $N$ is the number of nodes, $d$ is the dimensionality of node embedding, and $|\mathcal{T}_{v}|$ is the number of node types. Additionally, for the process of dynamic hypergraph learning, we use the incidence matrix H and nodes embedding X to update hyperedges embedding, with the computational cost of $\emph{O}\left(dMN\right)$. Note the number of hyperedges $M$ is proportional to the product of the $N$ and the number of edge types $|\mathcal{T}_{e}|$, thus the computational cost of hyperedges embedding updating is $\emph{O}\left(d|\mathcal{T}_{e}|N^{2}\right)$. Moreover, the computational cost of multi-head attention node embedding updating is $\emph{O}\left(d^{2}|\mathcal{T}_{e}|N\right)$, where $\mathcal{T}_{e}$ is the number of hyperedge types. Hence, the total computational cost of our proposal is $\emph{O}\left(d|\mathcal{T}_{e}|N^{2}+\left(d|\mathcal{T}_{v}|+d^{2}|\mathcal{T}_{e}|\right)N\right)$. The scale of $d$, $\mathcal{T}_{e}$ and $\mathcal{T}_{v}$ are usually far less than $N$, This computational cost is on par with many existing hypergraph-based models, such as [12] and [43].
4.5.2 Discussion
@@ 5 Experiment
  • RQ3: How much do the node embedding size and the number of multi-head $H$ affect the final results?
@@
5.1 Experimental Setup
Table 2: Statistics of the datasets.
@@ 5.1 Experimental Setup
5.1.1 Dataset
The experiments are conducted on three datasets, which are popularly used in classification tasks. The details of these datasets are summarized in Table 2.
@@ 5.1.1 Dataset
Table 3: Shared parameter setup.
@@
5.1.2 Baseline Models and Configurations
@@
The shared parameter setup of the framework can be found in Table 3. We randomly split all datasets into train/validate/test with the ratio of 0.2/0.1/0.7, respectively. It is worth mentioning that the hyperparameters of all baseline models are set to the optimal values reported in their original papers. All models are trained with a fixed 100 epochs, using an early stopping strategy when the performance on the validation set is not improved for 30 consecutive epochs. All trainable parameters of the neural network are initialized through Xavier [46] and optimized using Adam [45] with the learning rate 2e-3. The dropout rate is set to be 0.3 [47]. The optimal selection of hyperparameters of our model is further analysed in the following section.
Table 4: Performance of different models on node classification task.
@@
Figure 4: Comparison of F1 on IMDB_ALL.
5.2 Performance Analysis (RQ1)
We evaluate the performance of our proposed framework on node classification and link prediction tasks. Experiments for both tasks are repeated three times, and the average metric is reported.
5.2.1 Node Classification (RQ1.1)
In the node classification task, we initially perform a random sampling of nodes at varying percentages (10%, 30%, 50%, 70%) of the nodes from the three datasets (ACM, IMDB and DBLP). Then, any edges connected to nodes that are not sampled are eliminated. Subsequently, any sampled nodes that lack neighboring nodes are also removed. It is observed that a lower data sampling ratio correlates with an increased introduction of noise into the training dataset, due to the loss of original information. The best F1 results among all datasets are demonstrated in Table 4, in which the highest scores are marked in bold. We compared three different categories of graph learning models: homogeneous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraph learning models. Our proposed LFH achieves performance gain over other baseline models in all datasets by 2%-38%. Our model outperforms all baseline models in all datasets by an average improvement of 12.9%. Compared to four homogeneous pairwise GNN baselines GCN, GAT, GraphSAGE and GraphSAINT, the average F1 score improvement of LFH for all datasets is 14.4%. Regarding the cherry-picked heterogeneous pairwise graph learning models and hypergraph learning models, our model achieved an average improvement of 9.7% and 14.6%, respectively, compared to the best-recorded results in all datasets. It is worth noting that HGNN+ [20] is the only hypergraph learning baseline that utilizes pairwise information for hyperedge generation, resulting in better performance compared to other hypergraph learning models, except for our model. Furthermore, the training process of our model, which incorporates type-specific hyperedge generation and attentive embedding updates that exploits the heterogeneity of the graph, enables LFH to improve HGNN+ by an average of 5.7% in all datasets. Additionally, compared to the latest reporting heterogeneous learning model PC-HGN [8], the best results of our model gains an average increase by 5.2%. In general, we observe that our model outperforms all baselines for different data sample ratios (DSR), which is shown in Table 4. When the data sample ratios are set to be 70%, 50%, 30% and 10%, our model outperforms other models by 10.1%, 10.2%, 10.9% and 16.7% on average, respectively. It is worth noting that our model achieves higher average improvements when the data sampling ratio is set to 10%, and then the margin of average improvements declines as the data sampling ratio increases. To further verify the scalability of LFH, we conduct the experiment on the intact IMDB dataset, named IMDB_ALL, which contains 18405 nodes and 33973 edges. The experimental results on the IMDB_ALL demonstrate that LFH maintains superior performance over all other baselines, as shown in Fig. 4, with an average performance gain of around 7.3%.
Table 5: Performance of different models on link prediction.